OpenStudy (poohbear):

Perform the indicated operations and simplify the results. √8 (√2 + √3)

OpenStudy (anonymous):

\[√8 (√2 + √3)\rightarrow \sqrt{16}+\sqrt{24}\rightarrow 4+2\sqrt{6}\]